OpenStudy (anonymous):

graph 6x^2 +11x - 35

OpenStudy (anonymous):

can anyone help

OpenStudy (anonymous):

??

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (mrnood):

this is a quadratic equation - its graph is a parabola if you want to sketch it then if the value in front of x is positive then it is U shaped, if negative then it is inverted U If you set x = 0 in the equatin - that will give oyu the vlue where it crosses the Y axis if you solve the quadrtic =0 then that will give you the place where it crosses the X axis that is enough to sketch it

OpenStudy (anonymous):

6x^2 + 11x -35 = 0 Factor. To do this, Multiply -35 and 6 and get -210, Now what two numbers multiply to -210 and add up to 11 (the coefficient of the middle term)? Answer 21 and -10 So break the 11x into two terms. 6x^2 + 21x -10x -35 = 0 Group (6x^2 + 21x) + (-10x -35) = 0 Factor like terms 3x(2x + 7) + -5(2x + 7) = 0 Then factor out the (2x + 7) for (2x + 7)(3x - 5) = 0 so set each term = 0 and get x = 5/3 or x = -7/2

OpenStudy (anonymous):

6x² + 11x - 35 = 0 6x² - 10x + 21x - 35 = 0 2x(3x - 5) + 7(3x - 5) = 0 (3x - 5) (2x + 7) = 0 3x - 5 = 0 OR 2x + 7 = 0 x = 5/3 or x = -7/2

OpenStudy (anonymous):

6x^2 + 11x -35 = 0 6x^2 -10x + 21x -35 = 0 2x(3x-5)+7(3x-5)=0 (2x+7)(3x-5)=0 2x=-7 or 3x=5 x=-7/2 or x=5/3